OpenStudy (anonymous):

I need help solving this three variable equation 10x+6y+8z=100 10x+12y+6z=100 5x+4y+12z=100

OpenStudy (anonymous):

from eqn 1 and eqn 2, 10x+6y+8z=100=10x+12y+6z 10x+6y+8z=10x+12y+6z 6y+8z=12y+6z 3y+4z=6y+3z z=3y---------------eqn 4 from eqn 1 and eqn3 10x+6y+8z=100 =5x+4y+12z 10x+6y+8z=5x+4y+12z 5x+2y=4z now from eqn4 5x+2y=4(3y) 5x=10y x=2y--------------eqn 5 substituting value from eqn 4 and eqn 5 to eqn 1 10(2y)+6y+8(3y)=100 20y+6y+24y=100 50y=100 y=2 then x = 4 and z = 6